Phosphorus pentachloride decomposes according to the chemical equation PCl5(g)↽−−⇀PCl3(g)+Cl2(g)?c=1.80 at 250° C A 0.268 mol sample of PCl5(g) is injected into an empty 3.70 L reaction vessel held at 250° C. Calculate the concentrations of PCl5(g) and PCl3(g) at equilibrium.
Answer #1

molarity of PCl5 = 0.268 / 3.70 = 0.0724 M

PCl5 ---------------------------> PCl3 + Cl2

0.0724                                  0             0

-x                                         x              x

0.0724-x                             x              x

Kc = [PCl3][Cl2]/[PCl5]

1.80 = x^2 / 0.0724 -x

x^2 + 1.80 x - 0.1304 = 0

x = 0.0697

at equilibrium :

concentrations of PCl5   = 0.00270 M

concentrations of PCl3   = 0.0697 M